OpenStudy (anonymous):

Show your work. Solve for r. 5r-3(2r+5)=9

OpenStudy (anonymous):

5r-6r-15 = 9 -r = 24 r = -24

OpenStudy (anonymous):

why are ppl so against giving medals...they want the help....give the medal dammit